# Always-convex harmonic shears

### v1

@@ 2.1. Convex analytic functions onto half-planes or strips
where $a$ and $b$ are real numbers with $a<0<b$ and $b\leq|a|$.
@@
for some $K\in{\mathbb{C}}$ and some $\lambda$ of modulus one and different from $1$ and $-1$. This readily implies that
@@
Another application of Corollary 1 is that if the vertical shears of $\varphi$ are always-convex and $\varphi$ maps the unit disk onto a strip, then, necessarily, the parameter $\xi$ in (11) is either $-1$ or $1$, which implies that $\varphi\in\mathcal{A}$.
The previous examples are just a sample of the following well-known result proved, for instance, in [22, Thm. 2]: if an analytic function $\varphi$ maps the disk onto a convex domain, then $\varphi$ has a continuous extension to the closed unit disk (if $\infty$ is allowed as a value) and assumes no finite value more than once.
@@ 2.2. Analytic functions of bounded boundary rotation
[…]
@@
where $\mu(t)$ is a real-valued function of bounded variation on $[0,2\pi]$ with
[…]
The functions in $V_{k}$ are precisely those locally univalent analytic functions $\varphi$ normalized as in the class $\mathcal{S}$ which map ${\mathbb{D}}$ onto a domain of boundary rotation at most $k\pi$. We refer the reader to [19, 20] for the basic properties of the functions in $V_{k}$. We mention that if $k_{1}\leq k_{2}$, then $V_{k_{1}}\subset V_{k_{2}}$.
@@
Brannan proves the following (see [3, Thm. 3.9]).
@@
By adapting the proof of [3, Thm. 3.9], we can prove the following more general version of Theorem C.
@@
Notice that for $z=re^{i\theta}$ and $10\leq r<$ 1 we have
[…]
@@
[…]
@@
To finish this section, we mention the following theorem that will be fundamental for establishing our main result [18, Thm. 2.1].
@@
$E_{1}=\{e^{i\theta_{1}},e^{i\theta_{1}},\ldots,e^{i\theta_{n}}\}$
@@
[…]
@@ 3.1. Some useful results on convex harmonic mappings
It is known that if $f=h+\overline{g}\in S_{H}^{0}$ has constant boundary values, then $f$ maps the unit disk onto a polygon and its dilatation is a finite Blaschke product. This was proved by Sheil-Small in [26]. Hengartner and Schober established the converse. More precisely, they proved the following “surprising result” as qualified by themselves in [15, Thm.  3.3].
@@
The following theorem states some of the results from [1, Thm. 2.4].
@@
      $\lim_{t\to\theta^{-}}\widehat{f}(e^{it})\quad\text{to}\quad\lim_{t\to\theta^{+}}\widehat{f}(e^{it})$
@@ 3.2. The main theorem
=$\mathbb{H}=\{z\in{\mathbb{C}}\colon{\rm Re\,}z>-a\}$
@@
where $a$ and $b$ are real numbers with $a<0<b$ and $b\leq|a|$. Then, as is shown in [12], we can write, for some $A>0$ and some $\gamma\in[0,\pi/2)$,
$\mathbb{S}=\{z\in{\mathbb{C}}\colon A\left(\gamma-\frac{\pi}{2}\right)<{\rm Re\,}z<A\left(\gamma+\frac{\pi}{2}\right)\}\,.$
@@
- If $\varphi$ is bounded, then ${\rm Re\,}\varphi(e^{it})$ is continuous for all $t\in[0,2\pi)$.
@@
Case 3.1. Suppose that $E$ is empty. In this case, $f(e^{it})$ is constant, say $x_{0}+iy_{0}$, on $\partial{\mathbb{D}}$.
@@
Case 3.2. Let $E=\{e^{i\theta_{1}}\}$ for some $0\leq\theta_{1}<2\pi$. In this case, $f(e^{it})$ is equal to a constant, say $x_{1}+iy_{1}$, on $[0,2\pi)\setminus\{\theta_{1}\}$. As before, this implies that ${\rm Re\,}\varphi(e^{it})=x_{1}$ for all such $t$. The continuity of the boundary values of convex analytic functions again gives that $\varphi$ maps the unit disk onto a half-plane, so that $\varphi\in\mathcal{A}$.
Case 3.3. Let $E=\{e^{i\theta_{1}},e^{i\theta_{2}}\}$ for some $0\leq\theta_{1}<\theta_{2}<2\pi$. In this case, $f(e^{it})$ is equal to a constant, say $x_{1}+iy_{1}$, on $I_{1}=(\theta_{1},\theta_{2})$ and $x_{2}+iy_{2}$ on $t\in I_{2}=(0,2\pi)\setminus[\theta_{1},\theta_{2}]$. The continuity of ${\rm Re\,}\varphi(e^{it})$ at all points on the unit circle except perhaps for one gives $x_{1}=x_{2}$. Hence, the boundary of $\varphi({\mathbb{D}})$ is contained in a vertical line. That is, again $\varphi({\mathbb{D}})$ is a half-plane and $\varphi\in\mathcal{A}$.
Case 3.4. Finally, let $E=\{e^{i\theta_{1}},e^{i\theta_{2}},e^{i\theta_{3}}\}$ for some $0\leq\theta_{1}<\theta_{2}<\theta_{3}<2\pi$. In this case, $f(e^{it})$ is equal to a constant, say $x_{1}+iy_{1}$ on $I_{1}=(\theta_{1},\theta_{2})$, $x_{2}+iy_{2}$ on $I_{2}=(\theta_{2},\theta_{3})$, and $x_{3}+iy_{3}$ on $(0,2\pi)\setminus[\overline{I_{1}}\cup\overline{I_{2}}]$.
This gives that the boundary values ${\rm Re\,}\varphi(e^{it})$ are given by
@@
Another application of the fact that ${\rm Re\,}\varphi(e^{it})$ is continuous at all points on the unit circle except perhaps for one gives $x_{1}=x_{2}=x_{3}$, which allows us to conclude, once more, that $\varphi\in\mathcal{A}$.
Before continuying, it is convenient to notice that the choice of $\omega(z)=z$ in Case 3 could have been changed by any other dilatation of the form $\omega(z)=\lambda z^{N}$, where $|\lambda|=1$ and $N$ is a positive integer.
@@
The key tools used to resolve this case are Theorem C, Theorem D, and Theorem G.

### v2

@@ 2.1. Convex analytic functions onto half-planes or strips
where $a$ and $b$ are real numbers with $a<0<b$ and $|a|<b$.
@@
for some $K\in{\mathbb{C}}$ and some $\lambda$ of modulus one and different from $1$ and 1-1. This readily implies that
@@
Another application of Corollary 1 is that if the vertical shears of $\varphi$ are always-convex and $\varphi$ maps the unit disk onto a strip, then, necessarily, the parameter $\xi$ in (9) is either 1-1 or $1$, which implies that $\varphi\in\mathcal{A}$.
The previous examples are just a sample of the following well-known result proved, for instance, in [24, Thm. 2]: if an analytic function $\varphi$ maps the disk onto a convex domain, then $\varphi$ has a continuous extension to the closed unit disk (if $\infty$ is allowed as a value) and assumes no finite value more than once.
@@ 2.2. Analytic functions of bounded boundary rotation
$\int_{0}^{2\pi}\left|{\rm Re}\left(1+\frac{z\varphi^{\prime\prime}(z)}{\varphi^{\prime}(z)}\right)\right|\,d\theta\leq k\pi$
@@
where $\mu(t)$ is a real-valued function of bounded variation on $0,2\pi]$ with
$\int_{0}^{2\pi}d\mu(t)=2\pi\quad\text{and}\quad\int_{0}^{2\pi}|d\mu(t)|\leq k\pi\,.$
The functions in $V_{k}$ are precisely those locally univalent analytic functions $\varphi$ normalized as in the class $\mathcal{S}$ which map $\mathbb{D}$} onto a domain of boundary rotation at most $k\pi$. We refer the reader to [21, 22] for the basic properties of the functions in $V_{k}$. We mention that if $k_{1}\leq k_{2}$, then $V_{k_{1}}\subset V_{k_{2}}$. We refer the reader to [1], where analytic functions of bounded boundary rotation are used to generate close-to-convex harmonic mappings.
@@
Brannan proves the following (see [4, Thm. 3.9]).
@@
By adapting the proof of [4, Thm. 3.9], we can prove the following more general version of Theorem C.
@@
Notice that for $z=re^{i\theta}$ and $0\leq r<1$ we have
$\int_{0}^{2\pi}$ $\left|{\rm Re}\left(\frac{2\lambda z^{N}}{1-\lambda^{2}z^{2N}}\right)\right|\,d\theta=\frac{1}{2}\int_{0}^{2\pi}\left|{\rm Re}\left(\frac{1+\lambda z^{N}}{1-\lambda z^{N}}-\frac{1-\lambda z^{N}}{1+\lambda z^{N}}\right)\right|\,d\theta$
$\leq\frac{1}{2}\int_{0}^{2\pi}{\rm Re}\left(\frac{1+\lambda z^{N}}{1-\lambda z^{N}}\right)\,d\theta+\frac{1}{2}\int_{0}^{2\pi}{\rm Re}\left(\frac{1-\lambda z^{N}}{1+\lambda z^{N}}\right)\,d\theta=2\pi N\,.$
@@
$\int_{0}^{2\pi}\left|{\rm Re}\left(1+\frac{z\psi^{\prime\prime}(z)}{\psi^{\prime}(z)}\right)\right|\,d\theta\leq(k+2N)\pi.$
@@
To finish this section, we mention the following theorem that will be fundamental for establishing our main result [20, Thm. 2.1].
@@
$E_{1}=\{e^{i\theta_{1}},\ldots,e^{i\theta_{n}}\}$
@@
    $\frac{d\varphi(e^{i\theta})}{d\theta}=ie^{i\theta}\varphi^{\prime}(e^{i\theta})\quad\text{a.e.}\,$,
@@ 3.1. Some useful results on convex harmonic mappings
It is known that if $f=h+\overline{g}\in S_{H}^{0}$ has constant boundary values, then $f$ maps the unit disk onto a polygon and its dilatation is a finite Blaschke product. This was proved by Sheil-Small in [30]. Hengartner and Schober established the converse. More precisely, they proved the following “surprising result” as qualified by themselves in [16, Thm.  3.3].
@@
The following theorem states some of the results from [2, Thm. 2.4].
@@
      $\lim_{t\to\theta^{-}}\widehat{f}(e^{it})\quad\text{to}\quad\lim_{t\to\theta^{+}}\widehat{f}(e^{it})\,.$
@@ 3.2. The main theorem
$\mathbb{H}=\{z\in{\mathbb{C}}\colon{\rm Re\,}z>-a\}$
@@
where $a$ and $b$ are real numbers with $a<0<b$ and $|a|<b$. Then, as is shown in [13], we can write, for some $A>0$ and some $\gamma\in[0,\pi/2)$,
$\mathbb{S}=\left\{z\in{\mathbb{C}}\colon A\left(\gamma-\frac{\pi}{2}\right)<{\rm Re\,}z<A\left(\gamma+\frac{\pi}{2}\right)\right\}\,.$
@@
- If $\varphi$ is bounded, then ${\rm Re\,}\varphi(e^{it})$ is continuous for all $t\in[0,2\pi]$.
@@
Case 3.1. Suppose that $E$ is empty. In this case, $f(e^{it})$ is constant, say $x_{1}+iy_{1}$, on $\partial\mathbb{D}$.
@@
Case 3.2. Let $E=\{e^{i\theta_{1}}\}$ for some $0\leq\theta_{1}<2\pi$. In this case, $f(e^{it})$ is equal to a constant, say $x_{1}+iy_{1}$ (which is not necessarily equal to the constant in the previous case), on $0,2\pi)\setminus\{\theta_{1}\}$. As before, this implies that ${\rm Re\,}\varphi(e^{it})=x_{1}$ for all such $t$. The continuity of the boundary values of convex analytic functions again gives that $\varphi$ maps the unit disk onto a half-plane, so that $\varphi\in\mathcal{A}$.
Case 3.3. We now jointly consider the cases when either $E=\{e^{i\theta_{1}},e^{i\theta_{2}}\}$ for some $0\leq\theta_{1}<\theta_{2}<2\pi$ or $E=\{e^{i\theta_{1}},e^{i\theta_{2}},e^{i\theta_{3}}\}$ for some $0\leq\theta_{1}<\theta_{2}<\theta_{3}<2\pi$. The argument is similar in both cases and the conclusion is obtained as a consequence of the fact that ${\rm Re\,}\varphi(e^{it})$ turns out to be constant at all points on the unit circle except perhaps one, so that the result follows as before. In fact, the case when $E$ consists of $2$ points can be reduced to the one with $3$ points by setting $e^{i\theta_{2}}=e^{i\theta_{3}}$.
Let us then suppose that $f(e^{it})$ is equal to a constant, say $x_{1}+iy_{1}$ on $I_{1}=(\theta_{1},\theta_{2})$, $x_{2}+iy_{2}$ on $I_{2}=(\theta_{2},\theta_{3})$, and $x_{3}+iy_{3}$ on $(0,2\pi)\setminus[\overline{I_{1}}\cup\overline{I_{2}}]$.
This gives that the boundary values ${\rm Re\,}\varphi(e^{it})={\rm Re\,}f(e^{it})$ are given by
@@
Using again that ${\rm Re\,}\varphi(e^{it})$ is continuous at all points on the unit circle except perhaps for one gives $x_{1}=x_{2}=x_{3}$, which allows us to conclude, as announced, that $\varphi\in\mathcal{A}$.
Before continuing, it is convenient to notice that the choice of $\omega(z)=z$ in Case 3 could have been changed by any other dilatation of the form $\omega(z)=\lambda z^{N}$, where $|\lambda|=1$ and $N$ is a positive integer.
@@
The key tools used to resolve this case are Theorems  C, D, and G.
@@
Acknowledgements
M. Wołoszkiewicz-Cyll would like to thank the Department of Mathematical Analysis at the University of La Laguna for their hospitality during her research visit.
The authors would like to thank the referees for their careful reading of the manuscript, for their useful suggestions to improve the exposition, and for providing fundamental references for the bibliography. They are also grateful to Professors M. D. Contreras and L. Rodríguez-Piazza for bringing to their attention an inaccuracy in the first paragraph of an earlier version of the manuscript, which also appears in the published version but does not affect the results. This has been corrected in the current arXiv submission.